### v1

# Do Sell-side Analyst Reports Have Investment Value?

Abstract
This paper documents new investment value in analyst reports. Analyst narratives embedded with large language models strongly forecast future stock returns, generating significant alpha beyond established analyst-based and fundamental-based factors. The return predictability arises primarily from reports that convey negative sentiment but forecast favorable long-term prospects, suggesting systematic market overreaction to near-term negative news. The effect is more pronounced for large, mature firms and for reports authored by skilled, experienced analysts. A Shapley value decomposition reveals that analysts’ strategic outlook contributes the most to portfolio performance, especially forward-looking discussions on fundamentals. Beyond demonstrating untapped value in qualitative information, this paper illustrates the broader potential of artificial intelligence to augment—rather than replace—expert human judgment in financial markets.
JEL Classification: G12, G14, G24
Keywords: Analyst Report, Investment Value, Large Language Models, Explainable Artificial Intelligence
1 Introduction
Can narratives in sell-side analyst reports predict stock returns? While analysts provide both quantitative forecasts (earnings estimates, target prices, and recommendations) and qualitative analyses in written reports, decades of research have focused primarily on the former. Studies consistently show that quantitative forecasts generate profitable trading signals (e.g., Womack, 1996; Brav and Lehavy, 2003; Jegadeesh et al., 2004; Birru et al., 2022; Farago et al., 2023). Yet the narrative components of analyst reports remain an underutilized information source that may contain significant investment value.
The main challenge in analyzing report content has been the unstructured and high-dimensional nature of text data. Traditional approaches often condense text data to a single sentiment measure, but this dimensional reduction raises two critical concerns. First, when sentiment measures diverge from numerical forecasts, it becomes difficult to distinguish whether these gaps represent measurement error or genuine analyst insight. Second, dictionary-based sentiment methods face fundamental challenges in context-dependent interpretation, as highlighted by Loughran and McDonald (2011). The recent emergence of large language models (LLMs) offers a promising solution to these challenges by enabling more nuanced and comprehensive analysis of textual information.
In this paper, I analyze 1.2 million analyst reports from Thomson Reuters Investext between 2000 and 2023 to address two fundamental questions: whether analyst reports contain incremental investment value, and if so, what drives this value? Using this comprehensive dataset, I propose an integrated framework that examines investment value through three lenses: information representation, information relevance, and information interpretation.
First, I map analyst reports into a high-dimensional vector space using LLM transformer embeddings. These embeddings capture the LLM’s deep linguistic understanding and semantic interpretation of the text, representing a sophisticated “thinking process” that converts textual input into structured output.  [^1] This approach of text representation has gained traction in recent studies (Jha et al., 2020; Chen et al., 2022; Chen et al., 2023b). Second, I use machine learning models to extract value-relevant information to predict future 12-month stock returns and construct trading strategies based on these predictions. Finally, I conduct attribution analysis using Shapley value decomposition to identify which report components contribute most significantly to portfolio performance.
I begin by examining the return predictability of analyst reports with Ridge regressions. Projecting future 12-month stock returns on embeddings generated by the LLaMA3-8B model reveals significant out-of-sample predictive power. This predictability remains robust when controlling for quantitative analyst forecasts (recommendation revisions, earnings forecast revisions, and target price revisions) and when using DGTW-adjusted returns. Cross-sectionally, the predictive power is strongest for large, mature firms and reports authored by skilled, experienced analysts. This pattern suggests that text embeddings capture unique, longer-horizon insights that even well-covered stocks fail to fully price, and that superior information production requires both analytical capability and deep institutional knowledge.
To assess economic significance, I construct a monthly rebalanced portfolio sorted on analyst report return predictions. A self-financing trading strategy that buys stocks with the most positive predictions and shorts those with the most negative predictions generates an average monthly return of 1.04% with 68 basis points alpha (t = 2.64) relative to the Fama and French (2015) five factors plus momentum. To further evaluate the incremental value, I compare the strategy’s performance to 94 fundamental-based factors from Gu et al. (2020) and 18 analyst-based factors from Chen and Zimmermann (2022). The report-based strategy consistently generates significant information ratios ranging from 0.73 to 1.41, suggesting that analyst narratives contain substantial investment value beyond established predictive signals.
A natural question is whether analysts are simply piggybacking on known drift effects—such as the post-earnings-announcement-drift (PEAD) documented by Livnat and Mendenhall (2006)—and thus capturing documented price momentum rather than genuine new information. To investigate this, I examine how negative versus positive earnings shocks (SUE) interact with analysts’ predicted returns. Strikingly, most return predictability arises from reports predicting positive future returns but issued after negative news, suggesting a pattern distinct from simple price drift hypotheses.
This result raises a key question about where the market inefficiency originates: Are analysts attempting to correct market overreaction to negative news, or are they themselves contributing to the overreaction? To differentiate, I examine the sentiment and recommendations of reports with the highest and lowest return predictions. The results reveal that reports predicting positive returns tend to accompany downgrades, negative text sentiment, and negative contemporaneous returns. This pattern complicates the source of market inefficiency. If analysts were simply correcting the market’s overreaction, one would expect their recommendations to go against the prevailing negative sentiment. Instead, the reports show mixed signals: official downgrades coexist with bullish commentary on long-run prospects. The contradiction between these signals likely reduces the market’s ability to fully extract the positive information. Consequently, analysts both fuel short-term negativity and provide valuable insight about future rebounds—a dual role that highlights the importance of dissecting analysts’ qualitative content rather than relying solely on quantitative recommendations.
Another interesting question is which specific topics within analyst reports primarily drive investment value. To address this, I rely on a Shapley value decomposition framework that attributes a report’s predictive power to the constituted content. The basic idea is straightforward: after splitting the report into different topics, I test how trading strategies based on various combinations of these topics perform. By exhaustively examining all possible topic combinations, I obtain a fair estimate of how much each topic contributes to the final outcome.
The first step involves ensuring that each topic classification is both exclusive and meaningful, while avoiding arbitrary assumptions. I leverage the language capabilities of the ChatGPT-4o model to identify a hierarchy of topics, ultimately consolidating them into five meta-categories: Company and Industry Overview, Financial Analysis, Strategic Outlook, Risk and Governance, and Additional Content. I then fine-tune a BERT model to classify all 52 million sentences into these categories. Once the sentences are labeled, I construct topic-specific embeddings and generate return forecasts for each topic. By examining how portfolio performance changes when each topic is included or excluded—i.e., across all combinations—I derive Shapley values that quantify how much each category contributes to the Sharpe ratio and return.
I find that the Strategic Outlook discussions, despite comprising only 15% of report content, generate the largest contribution to investment value, accounting for 41% of the portfolio Sharpe ratio. By contrast, Financial Analysis—the largest category by distribution—contributes only about 16% of the Sharpe ratio, suggesting that standard financial data are quickly priced in, leaving less scope for future return prediction. Meanwhile, Company and Industry Overview and Risk and Governance show moderate yet meaningful contributions. Additional Content adds little. Portfolio performance validates the Shapley value attribution analysis. A trading strategy based solely on strategic content generates a monthly return of 1.41% with a Sharpe ratio of 0.93, significantly outperforming other content categories, which yield more modest returns between 0.52% and 0.58% monthly with Sharpe ratios of 0.37 to 0.41.
Strategic Outlook includes both investment thesis and valuation discussions, offering forward-looking elements such as growth potential, valuation, and investment opportunities. For instance, consider this excerpt from a report with a positive future return prediction despite a downgrade: “Although the shares are an attractive long-term investment, we must caution that the severity of the current industry downturn increases the chance of further near-term earnings surprises.” Linking the attribution result with the overreaction story, the investment opportunity appears to arise when analysts identify strong long-term prospects that are temporarily obscured by near-term challenges. Supportively, when dividing the strategic discussion further into three dimensions of timeframe, sentiment, and focus, the Shapley value decomposition emphasizes the contribution of sentences related to positive long-term fundamentals.
A potential concern about return predictability using LLMs is look-ahead bias (Sarkar and Vafa, 2024). Specifically, pre-trained language models could have learned future stock performance information from their training data, potentially causing information leakage through the embedding-generating process. However, I argue the results are robust to this concern for three reasons. First, LLMs are fundamentally designed to optimize language understanding rather than financial prediction capabilities. In other words, the pre-training process focuses on capturing semantic meaning through next-token prediction, not predicting stock returns. Second, if LLMs were indeed learning and exploiting future market information, we would expect to see the strongest predictability with the Company and Industry Overview content, as it contains the most comprehensive stock-specific information. Instead, this category generates insignificant alpha. Third, I directly test for look-ahead bias by examining portfolio performance after each model’s knowledge cutoff date using four different LLMs: BERT, RoBERTa, LLaMA2, and LLaMA3, which were trained on data ending between December 2018 and March 2023. The earliest models (BERT and RoBERTa) provide over five years of true out-of-sample testing. If the predictability stemmed from information leakage, performance should deteriorate in these post-knowledge cutoff periods. Instead, the strategies generate even superior performances, yielding monthly long-short returns between 1.76% and 5.28% with Sharpe ratios ranging from 0.74 to 2.37.
The paper contributes to the literature on the value of information from brokerage analysts. Previous papers generally focus on the cross-sectional return predictability of analyst quantitative outputs, such as stock recommendations (Green, 2006; Christophe et al., 2010; Chen et al., 2023a), earnings forecasts (Lys and Sohn, 1990; Bordalo et al., 2019), target price forecasts (Brav and Lehavy, 2003; Farago et al., 2023) and short-term trade ideas (Birru et al., 2022). I add to this strand of literature by documenting significant investment value in analyst report narratives that is distinct from and incremental to their quantitative forecasts. Furthermore, a contemporaneous complementary work by Lv (2024) employs similar embedding and Shapley value decomposition analytical framework to study analyst report information content and value for strategic investors. In contrast, this paper focuses on the investment value of report information in a longer horizon that is not immediately incorporated in stock prices. These papers complement each other in providing a comprehensive understanding of the dynamics of how analyst research content is processed by markets and creates value for investors.
This paper also contributes to the extensive overreaction literature by uncovering a novel pattern in analyst texts: negative sentiment paired with an overall positive long-run outlook.  [^2] For example: De Long et al. (1990); Porta et al. (1997); Hong and Stein (1999); Gennaioli et al. (2016); Weber (2018). A related paper by Bordalo et al. (2019) examines long-term growth expectations (LTG) and finds that analysts overreact to short-term news, with return predictability arising as prices gradually correct these expectations. A major distinction between our findings lies in the asymmetry of overreaction. While the LTG factor is mainly driven by analysts being over-optimistic and overvaluing stocks in high-LTG portfolios, my findings reveal that analysts often exhibit cautionary hedge—through downgrades or negative tones—while simultaneously embedding optimistic forecasts about longer-term fundamentals in their written discussions. The mismatch between near-term pessimism and constructive long-term insights provides a textual signature of market overreaction, where analysts behave cautiously with conservative narratives and forecasts despite being aware of positive fundamentals. By detecting these subtle, dual-layer signals in analyst reports, this work extends prior studies of De Bondt and Thaler (1990) on analyst overreaction and clarifies how overreaction dynamics are expedited with analyst written reports.
Third, this study contributes to the expanding literature on text factors, which has predominantly extracted sentiment or topics from corporate disclosures, earnings-call transcripts, and news media using dictionary-based or topic-modeling approaches (e.g., Tetlock, 2007; Chen et al., 2022; Meursault et al., 2023; Bybee et al., 2023; Hirshleifer et al., 2023). By contrast, I deploy state-of-the-art LLM embeddings to the nuanced, forward-looking narratives in sell-side analyst reports. These embeddings not only capture richer semantic and contextual relations but also have an additive property. The latter allows me to isolate specific content categories—such as financial analysis, risk factors, and strategic outlook—and then quantify each category’s standalone contribution to investment value through Shapley value decomposition. This approach offers a general framework for interpreting granular insights and attributing text predictive power to interested components.
The rest of the paper is organized as follows. Section 2 describes the data and methodology, including the prediction models, topic classification approach, and Shapley value decomposition framework. Section 3 presents the empirical results. Section 4 concludes.
2 Data and Methodology
2.1 Data
I combine analyst data from two primary sources. The first source is Thomson One’s Investext database, where I download 1,194,330 analyst reports covering S&P 1500 constituent firms from 2000 to 2023. Table 1 summarizes the annual statistics of analyst report characteristics, brokerage coverage, and unique analysts. The second source is the I/B/E/S database, which contains analyst qualitative forecasts including announcement and revision data. To merge these datasets, I follow a two-step process. First, I match lead analysts in Investext reports to I/B/E/S Analyst IDs (AMASKCD) following Li et al. (2024). Second, I link individual reports to their corresponding EPS forecasts, price targets, and recommendations in I/B/E/S using a five-day window around the announcement date ([-2, +2] days) following Huang et al. (2014).
Following Gu et al. (2020), I incorporate an array of 94 firm-level predictive characteristics. This dataset captures the market activity of all firms listed on the New York Stock Exchange (NYSE), American Stock Exchange (AMEX), and NASDAQ from January 1957 through December 2023. The monthly individual stock returns are sourced from the Center for Research in Security Prices (CRSP) database. Table A1 describes the details of anomaly variables.
To identify the incremental signal relative to quantitative analyst information, I compile 18 factors from Chen and Zimmermann (2022) with significant investment values in analyst-produced information documented in the literature. These factors use analysts’ forecasts, recommendations, or coverage changes to derive signals that predict stock performance. Table A2 provides a detailed description of the factors.
2.2 Methodology
2.2.1 Prediction Models
In transforming the report information into structured representations, I input each report text into the LLaMA3-8B model and take the average of 32 transformer layer embeddings as text representation.
To identify value-relevant information, I employ Ridge regression models to predict the return of stocks in the next 12 months following the report release using report embeddings. I deliberately employ linear Ridge regression models to emphasize the inherent information value of analyst reports rather than giving credit to sophisticated machine learning models. The Ridge regression models with an expanding training window are given by:
$Ret_{it,12m}=\beta_{0}+\beta^{\prime}y_{ijt}^{\mathrm{AI}}+\epsilon_{ijt},$ (1)
[…]
where $Ret_{it,12m}$ is the next 12 months return of stick $i$ following reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ is the structured representation (4,096-dimension embeddings generated with the LLaMA3-8B model) of content from analyst report $j$ on stock $i$ released at day $t$. I use the first five-year data as the initial training sample. The out-of-sample test starts in 2005 and ends in 2023.
2.2.2 Topic Modeling
To identify information sources, I first conduct a topic modeling exercise to group the sentences in analyst reports into clusters based on their content. Ideally, I would like to group sentences into exclusive topic categories that capture the full spectrum of information in analyst reports while avoiding subjective classification biases. I leverage the natural language understanding capabilities of state-of-the-art LLMs to extract exclusive and meaningful topics.
The hierarchical clustering algorithm begins with the following prompt applied to a random sample of 100 analyst reports:
Prompt 1: Please read the provided text file of sell-side analyst reports carefully. What are high-level mutually exclusive topics covered in these reports? Make sure that each sentence from the text file can be assigned to one of the topics. Here is the report content: {text}.
This prompting approach identifies 16 distinct topics: Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. To ensure complete coverage, I add a residual category “None of Above,” bringing the total to 17 topics.
For analytical tractability, I consolidate these topics into five broad categories that represent the main dimensions of analyst research: 1) company and industry overview; 2) financial analysis; 3) strategic outlook; 4) risk and governance; 5) additional content. Table LABEL:tab:_topic provides detailed descriptions of these categories and their constituent topics. Figure A1 presents word clouds of the categories.
I classify all 52,350,385 sentences into specific topics using a fine-tuned BERT model. I choose BERT for this classification task because its encoder architecture is specifically suitable for sequence classification compared to generative models like GPT. To train the classifier, I fine-tune BERT using a manually labeled dataset of 17,028 sentences extracted from 100 randomly selected analyst reports. The model assigns each sentence to one of the 17 identified topics, which are then mapped to the five meta-categories.
2.2.3 Shapley Value Decomposition
A follow-up question is where the investment value of analyst reports comes from. I attribute the investment value of the full report to five categories defined in Section 2.2.2 using the Shapley Additive exPlanations (SHAP) methodology proposed by Lv (2024).
SHAP, originally developed by Shapley (1953) in cooperative game theory, provides a framework for fairly allocating contributions among multiple players. The method satisfies key properties of efficiency and additivity, ensuring that the sum of all contributions equals the total value. By treating each word as a player, I can attribute the investment value contribution of individual words and their corresponding topics.
I implement this framework by mapping tokens to players, topics to teams, and monthly investment performance to individual games. The Shapley value decomposition process aggregates tokens within their respective topics, evaluates each topic’s performance across all possible coalitions in each monthly period, and computes the final contribution as the average marginal contribution of each topic across all periods. Below I present the mathematical formulations.
In implementing the Shapley framework, I first aggregate token embeddings into topic-level representations. LLMs generate token-level embeddings, which I combine into a report-level representation through averaging:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (3)
where $e_{i}$ is the average embedding of token $i$ across all transformer layers, $N$ is the total number of words in the report. Since each token is assigned to a specific topic, I can decompose the report embedding into topic-level representations with the average embeddings of all tokens that fall into the topic:
$y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (4)
where $P$ is the number of topics, $y_{p}^{emb}$ is the embedding for topic $p$, $N_{p}$ is the number of tokens belonging to topic $p$, and $i_{p}$ is the index of tokens within topic $p$.
With topic representations, I can generate Ridge predictions for all possible topic combinations and evaluate their investment performance. I calculate investment value outcomes in terms of Sharpe ratios or portfolio returns for each combination.
With outcomes from all possible topic combinations, I can compute the Shapley value using the following formula:
$\varphi_{p}\left(SR\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[SR\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-SR\left(y_{s}^{\text{emb}}\right)\right],$ (5)
where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set. Note that $\binom{n}{k}$ is the binomial coefficient. $y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. For each combination of topic embeddings, I calculate the Sharpe ratio accordingly. [^3] Shapley value requires an input for $X$ when it is missing. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fully distributes the portfolio Sharpe ratio among topics, revealing their relative contribution to report investment value.
Referring to Lv (2024), a potential concern in the topic representation stage is that token embeddings inherently contain information from other tokens, a feature designed into the LLM architecture to enable context understanding. One solution is to process embeddings at the sentence level rather than the report level to ban inter-sentence communications. The token-weighted mean of sentence embeddings can serve as an alternative report representation:
$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (6)
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embedding of the $i$-th sentence.
While this approach theoretically sacrifices contextual information across the entire report, I demonstrate in empirical tests that the performance deterioration is minimal.
3 Empirical Results
3.1 Investment Value
In this section, I examine the investment value of analyst reports through multiple lenses. Section 3.1.1 analyzes return predictability and portfolio performance. Section 3.1.2 compares the embedding-based model predictions with traditional report sentiment measures. Section 3.1.3 studies the cross-sectional variation in investment value across firm characteristics and analyst attributes.
3.1.1 Return Predictability
I begin with a univariate event-day approach to examine the stock price impact of analyst reports. This analysis is predicated on the notion that stock prices adjust when analysts release new information reflecting changes in their prior beliefs. The presence of abnormal returns following report releases would suggest that the market does not immediately fully incorporate analyst information. The price dynamic would indicate the investment value of the information.
To investigate these price dynamics, I use the predicted 12-month ahead stock returns from regression 1 as a measure of analysts’ outlook. I sort reports into decile groups based on their predictions for each trading day. [^4] In the sample, a trading day contains an average of 167 reports. To calculate buy-and-hold abnormal returns, I follow the characteristics-based benchmark approach proposed by Daniel et al. (1997):
$CAR_{0,T}=\prod_{t=0}^{T}\left(1+R_{it}\right)-\prod_{t=0}^{T}\left(1+R_{it}^{DGTW}\right),$ (7)
where $R_{it}$ represents the raw return of stock $i$ on trading day $t$, $R_{it}^{DGTW}$ denotes the value-weighted return of stock $i$’s benchmark portfolio on day $t$, and $T$ indicates the end date of the accumulation period. 00 represents the first trading day when the market is open at or after a report is released.
Figure 1 compares the abnormal returns between reports with the most positive and most negative outlooks over a 252-trading-day horizon (one calendar year). Reports with the highest predicted returns exhibit a cumulative DGTW-adjusted abnormal return of 3.75% over the subsequent 252 trading days, while those with the lowest predictions show an abnormal return of -0.56% over the same horizon. The return spread between High and Low groups is statistically significant at the 1% level. The price dynamics following report releases reveal an asymmetric pattern between positive and negative outlooks. This asymmetry suggests two potential explanations: either analysts are more skilled at identifying stocks with positive potential, or investors are not fully incorporating analysts’ positive narratives into prices.
As a second test of the investment value of analyst outlook, I use the following report-level regression model to explore return predictability:
$RET_{it,12m}=\alpha_{t}+\beta^{\prime}x_{i,t}+\varepsilon_{i,t+12},$ (8)
where $RET_{it,12m}$ represents the realized 12-month ahead return of stock $i$. $x_{i,t}$ denotes analyst information from each report, consisting of four measures: $\widehat{RET_{12m}}$ is the predicted 12-month ahead return from Ridge regression (1); $REC_{rev}$ represents recommendation revision, calculated as the difference between the current report’s recommendation and the analyst’s last recommendation for the same stock in I/B/E/S; $EF_{rev}$ denotes earnings forecast revision; and $TP_{rev}$ represents target price revision. The model includes firm fixed effects $\alpha_{t}$ to control for systematic market conditions. Standard errors are two-way clustered at both stock and year-quarter levels.
I report the regression results in Table 2. The estimated returns derived from analyst report information demonstrate statistically significant predictive power for future stock returns. After controlling for analyst-year-month and industry-year-month fixed effects, the coefficient in column (2) (0.06, t-statistics=4.37) suggests that a one standard deviation increase in estimated returns corresponds to a 10.41% increase in realized returns relative to the sample mean over the following year. This finding aligns with the evidence in Figure 1, confirming the consistency between analyst report outlook and subsequent stock market movements.
I further compare the predictive power of numerical analyst forecasts at the one-year horizon. Columns (3)-(6) show that the predictive power of text-based estimates remains robust when controlling for analyst recommendation revisions, earnings forecast revisions, and target price revisions. All coefficients maintain their statistical and economic significance. While these numerical forecasts are documented as valuable predictors in the literature, their coefficient estimates become statistically insignificant when competing with text predictions for one-year returns. This indicates that the report text contains unique quantitative information in forecasting stock market returns.
I proceed with a portfolio analysis to evaluate the economic value of analyst report predictions. The trading strategy design incorporates two key elements. First, portfolios are rebalanced monthly to balance the tradeoff between capturing new information and minimizing transaction costs. Second, motivated by the persistence in return predictability documented in Figure 1, I aggregate analyst report predictions over extended historical windows to capture sustained components in analyst outlooks.
Specifically, at the beginning of each month, I construct portfolios using the following procedure: First, for each stock, I calculate the average predicted return from all analyst reports issued during the past $LB$ months, where $LB\in\{9,12,18,24\}$. I then sort stocks into decile portfolios based on these averaged predictions. Within each decile portfolio, stocks are weighted by their previous month-end market capitalization to mitigate the influence of microcaps. I then construct a zero-cost long-short portfolio that takes a long position in the highest predicted return decile (decile 10) and a short position in the lowest predicted return decile (decile 1).
Table 3 presents performance statistics of value-weighted decile portfolios formed based on Ridge models’ return forecasts using different lookback periods (LB) of 9, 12, 18, and 24 months. While the decile performance shows some variation, the highest decile consistently outperforms the lowest decile. The long-short portfolio (H-L) that buys the highest decile and shorts the lowest decile generates positive returns ranging from 0.87% to 1.16% per month, with corresponding Sharpe ratios between 0.62 and 0.69. The portfolio alphas relative to the Fama-French five-factor and momentum factor are positive and statistically significant, ranging from 0.52 to 0.79 (t-statistics between 1.99 and 2.64). The alpha across decile portfolios reveals that the long-short portfolio performance is primarily driven by the long leg rather than the short leg, consistent with the asymmetric pattern documented in Figure 1. For subsequent analyses, I adopt the 12-month lookback period as the benchmark specification, as it yields the highest t-statistic (2.64) for the long-short portfolio alpha and aligns with Ridge models’ 12-month return prediction target.
Figure 2 plots the cumulative returns over the period of January 2005 to June 2024 for the 12-month lookback period trading strategy. A self-financing value-weighted portfolio that buys stocks with the most positive outlook and shorts those with the most pessimistic outlook generates a cumulative return of 846% over this period. For comparison, the value-weighted market return in excess of the risk-free rate during this period is 399%. Notably, the long lag earns 2274% over the 20 years, while the short leg delivers a cumulative excess return of 151%. Thus, the investment value stems from picking stocks with strong future prospects, which is consistent with the portfolio statistics in Table 3. The strategy experiences significant drawdowns during the global financial crisis and again post-2022, conditions in which fundamental signals can be temporarily overshadowed by systemic risk aversion and forced liquidations.
3.1.2 Short-term Overreaction
Analysts often release reports in response to major news or corporate announcements. This raises the question of whether the text predictability stems from documented price drift phenomena, such as the post-earnings-announcement-drift (PEAD), rather than reflecting genuine new information about firms’ long-term fundamentals.
To distinguish between the two drivers, I conduct a double-sorting analysis based on earnings surprises and return predictions, presented in Figure 3. First, reports are classified into ‘SUE(Small)’ or ‘SUE(Large)’ groups based on whether their stocks’ most recent earnings surprise falls below or above the monthly sample median. Following Livnat and Mendenhall (2006), earnings surprise (SUE) is calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement, with higher values indicating positive news. Second, within each SUE group, I sort reports into deciles based on their predicted returns. Recall that the return predictability is driven by long lag (positive predictions), if the observed predictability were attributable to a price‐drift like PEAD, then reports with high SUE (i.e., good earnings news) should exhibit a greater spread between the ‘High’ and ‘Low’ groups.
Surprisingly, the data reveal the opposite pattern. As shown in Figure 3 panel (a), the return spread is 7.52% for reports issued after low earnings surprises, compared to 4.30% for those issued following high surprises. Thus, it appears that weak earnings announcements in fact contribute more to return predictability, contradicting a drift‐based explanation and posing its own puzzle.
To gather more evidence, I examine several measures across the return prediction deciles: (1) $CAR_{[0,+1]}$, the two-day DGTW characteristic-adjusted buy-and-hold abnormal returns following report release dates, which captures immediate market reactions; (2) $REC_{rev}$, analyst recommendation revisions that reflect changes in analyst opinions; and (3) two text sentiment measures generated using a fine-tuned BERT model that classifies each sentence as positive, negative, or neutral. $Tone_{Head}$ evaluates report headlines, while $Tone_{Body}$ assesses the report body content. Specifically, $Tone_{Body}$ is calculated as:
$Tone_{Body,i}=\frac{N_{i}^{+}-N_{i}^{-}}{N_{i}},$ (9)
where $N_{i}^{+}$ ($N_{i}^{-}$) represents the number of positive (negative) sentences in report $i$ as classified by the BERT model, and $N_{i}$ denotes the total number of sentences in the report.
Table 4 presents two noteworthy findings. First, the realized returns ($R_{12m}$) exhibit a significant spread of 5%, which closely aligns with the ’H-L’ return spread documented in Figure 1. Second, the contemporaneous market reaction ($CAR_{[0,+1]}$) following report releases shows a negative relationship with predicted returns, with a significant high-minus-low spread of -0.25%. Together with the evidence in Figure 3, these results present an intriguing puzzle: The most optimistic forward-looking reports tend to follow bad news, and vice versa. I propose three potential explanations for this pattern. First, both analysts and the market may overreact to short-term negative news (e.g., a weak quarter or a disappointing press release). Yet the analyst’s deeper analysis (e.g., the model-based 12-month forecast) might conclude that the fundamentals remain solid, eventually driving future price appreciation. Under this scenario, we should observe consistent negative relationships between analyst recommendations, narrative tones, and predicted returns.
An alternative story is that analysts intentionally address short-term negativity but emphasize the firm’s longer-term strengths in order to reassure overly pessimistic investors. Here, the analyst acknowledges current problems but stresses the company’s strategy or fundamentals as a counterpoint to the short-run overreaction. If this is the driving mechanism, there should be a positive correlation between the analyst’s optimistic return forecasts and the tone of the report text.
A third explanation posits that analysts systematically include disclaimers or cautious language—even when they maintain a constructive long-term view—to hedge official recommendations and manage liability. In this case, the overall textual tone might appear muted or slightly negative despite a fundamentally bullish numerical recommendation.
The empirical evidence points to systematic overreaction to near-term negative signals, rather than a brief misalignment or benign hedging of recommendations. Table 4 shows that the most optimistic long-horizon forecasts consistently coincide with negative short-run sentiment (in both the text and recommendation revisions). Meanwhile, Figure 3 indicates that this predictability is strongest specifically among reports issued after weak earnings news and downgrades, suggesting analysts are not offsetting short-term setbacks with unequivocally bullish outlooks.
If overreaction is the primary driver of return predictability, does this imply that analysts’ forward-looking outlooks are irrelevant? To explore this, I conduct an analysis by sorting portfolios based on sentiment measures to assess whether such strategies can generate alpha. Table A4 reveals that simple sentiment-based sorting fails to produce significant alpha, suggesting that sentiment alone does not represent the informational value embedded in analyst reports. Rather, analysts’ assessments of fundamental value remain crucial for predicting subsequent return reversals.
Taken together, the data reveal that analysts and the market place disproportionate weight on immediate setbacks, even while the analyst’s expertise anticipates subsequent stock outperformance. In other words, the investment value arises from analysts’ ability to recognize longer-term fundamentals despite their own cautious or negative language in the short run. This systematic mismatch—where near-term negativity coincides with strong future returns—ultimately generates the portfolio alpha.
3.1.3 Cross-sectional Analyses
I next turn to the question of reports from what kinds of analysts on which stocks generate investment value. Figure 4 presents DGTW-adjusted buy-and-hold abnormal returns sorted on analyst and firm characteristics. The results reveal substantial heterogeneity in the investment value of research reports across different segments of the market.
I first examine the return predictability of analyst reports across large vs small stocks and mature vs young stocks. Panel (a) shows that reports on large-capitalization stocks generate significantly higher abnormal return spreads compared to small-cap stocks (4.14% vs 1.74%), with the difference particularly pronounced for stocks predicted to have high returns. This pattern suggests that despite the greater analyst coverage and presumed market efficiency in the large-cap segment, skilled analysts can still identify meaningful mispricing opportunities among these stocks. The life cycle results in Panel (b) indicate that mature firms offer greater potential for profitable analyst recommendations compared to young firms. Reports on mature firms predicted to have high returns generate cumulative abnormal returns of approximately 5.15% over the 252-day horizon, while similar reports on young firms yield a substantially lower return spread of 1.51%.
In addition, I also test the predictability in the cross-section of analyst skills and experience. Analyst skill, measured by historical EPS forecast accuracy and shown in Panel (c), appears to be a crucial determinant of report value. High-skill analysts generate significantly larger abnormal return spreads (5.74%), particularly for their most optimistic recommendations, while reports from low-skill analysts show significantly smaller investment value in terms of return spreads (4.44%). Finally, Panel (d) demonstrates that analyst experience plays a vital role in report effectiveness. Experienced analysts’ reports, especially those predicting high returns, generate substantial abnormal returns approaching 6.14% over the holding period. In contrast, reports from less experienced analysts show less ability to identify profitable investment opportunities.
These findings suggest that the investment value of research reports is concentrated among experienced, skilled analysts covering large, mature firms. This pattern is consistent with the idea that superior information production requires both analytical capability and deep institutional knowledge. Moreover, large, mature businesses with more predictable cash flows and stable competitive moats often snap back from negative news—leading to a positive return drift that skilled analysts can exploit.
3.2 What content matters?
I now investigate the sources of investment value in analyst reports by decomposing portfolio returns and Sharpe ratios across five topic categories, as described in Section 2.2.2. Using yearly fitted Ridge models, I calculate each topic’s contribution to portfolio performance by measuring the average decrease in performance across all possible topic combinations. I group together topic Shapley values with specific categories and report the sum of their constitute importance.
Table 5 presents both the distribution of content and its contribution to portfolio performance. Financial analysis and company/industry overview dominate the content, comprising 36.56% and 28.53% of total coverage respectively. Strategic outlook and risk and governance form a second tier of coverage at 15.13% and 14.14% of sentences, while additional content accounts for the remaining 5.63%.
The Shapley value decomposition reveals a different pattern. Strategic Outlook emerges as the most valuable category, accounting for 41.34% and 31.43% of the portfolio’s Sharpe ratio and returns respectively, despite not being the most extensively discussed component. Company and Industry Overview maintains proportional importance, contributing 27.61% to the Sharpe ratio and 26.92% to returns. Risk and Governance shows outsized impact on returns relative to its discussion frequency, accounting for 11.21% of Sharpe ratio value and 21.36% of returns. Additional Content has minimal impact on performance.
Perhaps the most surprising result is the low contribution of Financial Analysis given findings in Lv (2024) that emphasize high information content in income statement analysis. Despite constituting the most extensively covered category (36.56% of content), Financial analysis discussions contribute only to 16.39% of the Sharpe ratio and 19.53% of returns. One reason may be that financial data are more rapidly priced in, leaving less room for drift or mispricing to persist. In contrast, strategic and forward-looking insights may only gradually be reflected in valuations, thus creating greater opportunity for abnormal returns.
The findings in Table 5 suggest that the investment value of analyst reports stems primarily from their strategic insights and industry analysis rather than traditional financial analysis. The substantial contribution of strategic content to portfolio performance, despite its relatively lower coverage, indicates that analysts’ forward-looking assessments and strategic perspectives may be particularly valuable for investment decisions.
Table 6 presents a detailed analysis of strategic outlook content along three key dimensions - timeframe, sentiment, and analytical focus - using Shapley value decomposition. For each dimension, I classify sentences using a structured prompt that categorizes timeframe (long-term, short-term, or both), sentiment (negative, neutral, or positive), and focus (risk versus fundamental analysis) with the following prompt:
Prompt 2: Please read the following sentence from a sell-side analyst report (investment thesis section) carefully and classify it into three numeric labels: Timeframe: 1 = Long-term (multi-quarter or multi-year outlook); 2 = Short-term (next quarter or near-term); 3 = Both (if it clearly addresses both short- and long-term). Sentiment: 1 = Positive potential; 0 = Neutral; 1 = Negative/Bearish. Focus: 1 = Fundamentals (e.g., belief in growth strategy, industry tailwinds, earnings); 0 = Cautious risk (e.g., warnings, near-term headwinds, legal/regulatory risk). Important note: If the sentence mentions both short-term risk and a fundamental (longer-term) driver, prioritize the fundamental aspect and mark the third label as 1 (Fundamentals). Output format: Only output the three numbers in the format: [X, Y, Z]. Here is the sentence: {sentence}
The results reveal several notable patterns. First, long-term oriented discussions dominate the investment value, accounting for approximately half of both Sharpe ratio (49.51%) and return (50.20%) contributions. Short-term discussions contribute about a quarter of the predictive power, while discussions spanning both timeframes account for the remaining portion.
Second, the sentiment analysis shows that positive outlooks generate the largest share of investment value, contributing 49.94% to Sharpe ratio and 50.86% to returns. Neutral and negative sentiments have roughly equal contributions, each accounting for about a quarter of the portfolio performance metrics.
Perhaps most strikingly, the focus decomposition shows that fundamental analysis overwhelmingly drives investment value, accounting for 87.00% of Sharpe ratio and 87.75% of return contributions. Risk-focused discussions, while important for comprehensive analysis, contribute relatively little to portfolio performance metrics.
These findings suggest that the predictive power of strategic outlook sections stems primarily from long-term, positive discussions focused on fundamental analysis, rather than from risk warnings or short-term forecasts. This pattern aligns with the notion that durable competitive advantages and long-term growth opportunities require more analytical skill to identify and value correctly than short-term fluctuations or risk factors.
A follow-up question would be, since Shapley value attribution pins down the content contributing the most to investment value, can we separate this most valuable information and form a portfolio free of text noise? In Table 7, I report the value-weighted decile portfolio performance sorted by category-specific content. I focus on four categories, excluding additional content that primarily consists of appendices and latent sentences.
Strategic Outlook content generates the strongest portfolio performance, with the long-short strategy delivering a mean return of 1.41% per month and a Sharpe ratio of 0.93. The outperformance appears to be driven by both the long and short legs, with the highest decile earning 1.87% monthly and showing significant alpha (0.72, t = 2.85). This superior performance aligns with the earlier Shapley value decomposition, which identifies Strategic Outlook as the most valuable content category.
The remaining categories - Company and Industry Overview, Financial Analysis, and Risk and Governance - generate more modest performance metrics, with long-short returns ranging from 0.52% to 0.58% monthly and Sharpe ratios between 0.37 and 0.41. While the highest deciles of these categories show some predictive power, particularly Risk and Governance with an alpha of 0.32 (t = 1.97), their overall performance is substantially weaker than Strategic Outlook. This pattern is consistent with the Shapley value decomposition and suggests that traditional financial analysis, industry insights, and risk assessments are more efficiently priced by the market than forward-looking strategic insights.
Therefore, isolating Strategic Outlook content may indeed help form more efficient portfolios by focusing on the most valuable component of analyst research. Confirming the effectiveness of the attribution analysis with Shapley values, the results again highlight the long-term investment opportunities embedded in analysts’ forward-looking strategic insights.
3.3 Incremental Investment Value
Given the ”zoo of factors” documented to have investment value, a key question emerges: is the investment value of analyst reports incremental to existing predictive signals from analyst forecasts and fundamental factors? I therefore extend my analysis to examine whether analyst reports contain unique information beyond two comprehensive sets of established predictors: 18 analyst-based factors including EPS forecast revisions and recommendation changes from the Chen and Zimmermann (2022) database, and 94 market and fundamental-based factors documented to have cross-sectional predictability in Gu et al. (2020). [^5] The open-source code is publicly available on Dacheng Xiu’s website. Detailed descriptions of these analyst-based and fundamental-based factors can be found in Tables A2 and A1, respectively.
In Table 8, I evaluate the predictive power of analyst reports relative to and in combination with established return predictors. The analysis examines four main signals: analyst report full content predictions ($RP$), strategic outlook content predictions ($SO$), the average return series of 18 analyst-based factors ($ANA$), and predictions from a 4-layer artificial neural network ($ANOM$) - the best-performing machine learning model in Gu et al. (2020). For each strategy and combination, I report mean returns, Sharpe ratio (SR), and information ratio (IR), with corresponding t-statistics in parentheses. To assess the robustness of the predictive signals, I compute risk-adjusted returns ($\alpha$) relative to four prominent factor models: the Fama and French (2015) five factors ($\alpha_{F5}$), the five factors plus momentum ($\alpha_{F6}$), the Hou et al. (2015) q-factor model ($\alpha_{HXZ}$), and the Daniel et al. (2020) behavioral factor model ($\alpha_{DHS}$). The Information Ratio of the portfolio $i$ relative to a benchmark $b$ is calculated as:
$IR_{i,b}=\frac{\bar{r}_{i,b}}{\sigma_{i,b}},$ (10)
where $\bar{r}_{i,b}$ is the average active return (difference between portfolio $i$ and benchmark $b$ returns) and $\sigma_{i,b}$ is the standard deviation of these active returns, also known as tracking error. In Table 8, I compute the IR of combined strategies relative to individual components to measure the incremental value of information sources. For example, the IR of “RP + ANA” measures the incremental gain from combining report predictions with analyst-based factors, relative to using “ANA” signals only.
Panel A first establishes the standalone performance of each information source. Both full report content (RP) and strategic outlook sections (SO) generate significant risk-adjusted returns across various factor models, with alphas ranging from 0.73 to 1.21 for RP and 0.95 to 1.46 for SO. The traditional analyst-based factors (ANA) show modest but significant predictive power (alphas between 0.16 and 0.28). The machine learning-based anomaly portfolio (ANOM) exhibits strong performance with alphas between 1.15 and 1.40, suggesting effective capture of fundamental signals.
Panels B and C address the central question of incremental value. When combining report predictions with analyst-based factors (RP + ANA), the portfolio maintains significant alphas (0.50-0.68) and shows meaningful information gain with an IR of 0.73. The combination with fundamental factors (RP + ANOM) yields even stronger results, with alphas between 1.03 and 1.18 and an IR of 1.16. Most notably, incorporating both analyst and fundamental factors (RP + ANA + ANOM) produces the highest Sharpe ratio (1.60) and consistent alphas across factor models (0.77-0.92), with an IR of 1.23 (t=4.05). The patterns are similar but even stronger when using strategic outlook content instead of full reports (IR=1.41, t=4.35), consistent with my earlier findings about the particular value of strategic discussions.
These results suggest that analyst reports contain statistically and economically significant incremental information not captured by traditional analyst outputs like forecast revisions or by machine learning models trained on fundamental data. The strong incremental performance and information ratios indicate that textual analysis of research reports, particularly their strategic content, offers a distinct and complementary source of investment value.
3.4 Look ahead Bias
A key concern in demonstrating out-of-sample profitability from analyst reports is look-ahead bias. Because LLMs are pre-trained on massive text corpora, there is a risk that they could have indirectly learned from future stock performance data. While Table 7 shows that providing corporate and industry information to LLMs does not generate significant profits—mitigating concerns about entity- or industry-level information leakage—one might argue that strategic outlooks contain forward-looking information that could align with the LLMs’ pre-training corpus. The fact that these outlooks can later be confirmed by actual outcomes may introduce a potential source of information leakage.
To address this concern, I conduct a stringent analysis restricted to post-knowledge cutoff samples. This analysis encompasses a broad range of language models with diverse architectures and scales. BERT and RoBERTa are relatively smaller, pioneering transformer-based NLP models (with 110 million to 355 million parameters). In contrast, LLaMA2-13B is a significantly larger model at 13 billion parameters, while LLaMA3-8B represents a more recent architecture with 8 billion parameters. Specifically, I evaluate portfolio performance only after each model’s training data cutoff date: BERT (post-June 2024), RoBERTa (post-June 2024), LLaMA2 (post-June 2024), and LLaMA3 (post-April 2024). Table 9 presents these results across two dimensions: analysis of full reports (Panel A) and strategic outlooks (Panel B).
Strikingly, the performance in this true out-of-sample period is even stronger than in the full sample. Examining the high-minus-low (H-L) portfolio returns in Panel A, all models generate substantial positive returns ranging from 1.59% (LLaMA2) to 2.70% (LLaMA3) with corresponding Sharpe ratios between 0.74 and 1.28. The performance amplifies further when focusing solely on strategic outlooks (Panel B), with H-L returns more than doubling for some models, ranging from 1.81% (RoBERTa) to an impressive 5.28% (LLaMA3). The enhanced performance in the post-cutoff period provides compelling evidence that the LLMs’ ability to extract profitable information from analyst reports is not driven by potential information leakage from their training data.
4 Conclusion
Investment value from sell-side analyst forecasts has been well-documented in the literature. Beyond offering simple numerical projections, analysts also produce research reports that detail their fundamental analyses and valuation methods—reports that institutional investors often deem more valuable than forecasts alone. Building on this notion, I conduct a comprehensive examination of a large corpus of analyst reports spanning 2000 to 2023, extracting embedded value-relevant information using state-of-the-art LLMs and machine learning techniques.
My findings reveal a new source of investment value. The textual information, when represented as structured vectors, predicts future stock returns. This return predictability offers significant incremental investment value beyond existing analyst- and fundamental-based factors, and it remains robust in a period free from look-ahead bias.
I also observe that return predictability is predominantly driven by reports conveying negative sentiment yet predicting positive future returns. This pattern suggests a systematic overreaction to negative news—such as weak quarterly earnings—even when a stock’s fundamentals remain strong. Further, the predictability is more pronounced for large, mature stocks and is particularly evident in reports authored by skilled and experienced analysts.
Next, I investigate which specific content within analyst reports contributes most to their investment value. Using a Shapley value decomposition, I show that Strategic Outlook content—especially statements revealing positive future fundamentals in the investment thesis and valuation models—dominates portfolio performance. Remarkably, a trading strategy based solely on strategic outlook content outperforms trading strategies based on full content.
This study provides the first direct evidence of the investment value embedded in analyst report text. These reports reveal analysts’ beliefs and expertise regarding particular stocks and convey valuable information to the market. More broadly, the findings illustrate the potential of artificial intelligence to enhance human judgment in financial contexts. Rather than replacing analyst insights, machine learning adds the most value when scaling and systematically extracting human-generated information from natural language.
In addition, the Shapley value decomposition approach proves to be a powerful tool for interpreting textual information, helping researchers and investors understand how different content types affect outcomes of interest. Future research could extend these methods to other forms of financial narratives and examine the mechanisms through which qualitative information is gradually incorporated into asset prices.
Appendix
Appendix A Tables and Figures
(a) (1) Company and Industry Overview
(b) (2) Financial Analysis
(c) (3) Strategic Outlook
(d) (4) Risk and Governance
Figure A1: Word Clouds of Topics
This figure presents word clouds for 4 categories frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The categories include Company and Industry Overview, Financial Analysis, Strategic Outlook and Risk and Governance. The “Additional Content” category is excluded from the visualization.
Table A1: Anomaly Variables
This table reports the market and firm-related fundamental factors considered in this paper. I follow Gu et al. (2020) to construct the dataset. Gu et al. (2020) provide a detailed description of the dataset in their appendix.
Acronym Description Acronym Description
absacc Absolute accruals mom36m 36-month momentum
acc Working capital accruals mom6m 6-month momentum
aeavol Abnormal earnings announcement volume ms Financial statement score
age # years since first Compustat coverage mvel1 Size
agr Asset growth mve_ia Industry-adjusted size
baspread Bid-ask spread niincr Number of earnings increases
beta Beta operprof Operating profitability
betasq Beta squared orgcap Organizational capital
bm Book-to-market pchcapx_ia Industry adjusted % change in capital expenditures
bm_ia Industry-adjusted book to market pchcurrat % change in current ratio
cash Cash holdings pchdepr % change in depreciation
cashdebt Cash flow to debt pchgm_pchsale % change in gross margin - % change in sales
cashpr Cash productivity pchquick % change in quick ratio
cfp Cash flow to price ratio pchsale_pchinvt % change in sales - % change in inventory
cfp_ia Industry-adjusted cash flow to price ratio pchsale_pchrect % change in sales - % change in A/R
chatoia Industry-adjusted change in asset turnover pchsale_pchxsga % change in sales - % change in SG&A
chcsho Change in shares outstanding pchsaleinv % change in sales-to-inventory
chempia Industry-adjusted change in employees pctacc Percent accruals
chinv Change in inventory pricedelay Price delay
chmom Change in 6-month momentum ps Financial statements score
chpmia Industry-adjusted change in profit margin quick Quick ratio
chtx Change in tax expense rd R&D increase
cinvest Corporate investment rd_mve R&D to market capitalization
convind Convertible debt indicator rd_sale R&D to sales
currat Current ratio realestate Real estate holdings
depr Depreciation / PP&E retvol Return volatility
divi Dividend initiation roaq Return on assets
divo Dividend omission roavol Earnings volatility
dolvol Dollar trading volume roeq Return on equity
dy Dividend to price roic Return on invested capital
ear Earnings announcement return rsup Revenue surprise
egr Growth in common shareholder equity salecash Sales to cash
ep Earnings to price saleinv Sales to inventory
gma Gross profitability salerec Sales to receivables
grCAPX Growth in capital expenditures secured Secured debt
grltnoa Growth in long term net operating assets securedind Secured debt indicator
herf Industry sales concentration sgr Sales growth
hire Employee growth rate sin Sin stocks
idiovol Idiosyncratic return volatility sp Sales to price
ill Illiquidity std_dolvol Volatility of liquidity (dollar trading volume)
indmom Industry momentum std_turn Volatility of liquidity (share turnover)
invest Capital expenditures and inventory stdacc Accrual volatility
lev Leverage stdcf Cash flow volatility
lgr Growth in long-term debt tang Debt capacity/firm tangibility
maxret Maximum daily return tb Tax income to book income
mom12m 12-month momentum turn Share turnover
mom1m 1-month momentum zerotrade Zero trading days
Table A2: Analyst Variables
This table reports the analyst factors considered in this paper. I compile the data from Chen and Zimmermann (2022) website.
Acronym Journal (Publish Year) Description
AnalystRevision FAJ (1984) EPS forecast revision
AnalystValue JAE (1998) Analyst Value
AOP JAE (1998) Analyst Optimism
ChangeInRecommendation JF (2004) Change in recommendation
ChForecastAccrual RAS (2004) Change in Forecast and Accrual
ChNAnalyst ROF (2008) Decline in Analyst Coverage
ConsRecomm JF (2001) Consensus Recommendation
CredRatDG JF (2001) Credit Rating Downgrade
DownRecomm JF (2001) Down forecast EPS
EarningsForecastDisparity JFE (2011) Long-vs-short EPS forecasts
ExclExp RAS (2003) Excluded Expenses
FEPS WP (2006) Analyst earnings per share
fgr5yrLag JF (1996) Long-term EPS forecast
ForecastDispersion JF (2002) EPS Forecast Dispersion
Recomm_ShortInterest AR (2011) Analyst Recommendations and Short-Interest
REV6 JF (1996) Earnings forecast revisions
sfe AR (2001) Earnings Forecast to price
UpRecomm JF (2001) Up Forecast
Table A3: Topic Categories and Descriptions
Topic Desriptions
Company and Industry Overview
Executive Summary Provides a high-level overview of the report’s key findings and conclusions; includes a brief description of the company, its industry, and the purpose of the report; highlights the most important points from the analysis, such as the company’s financial performance, competitive position, and growth prospects.
Company Overview Offers a comprehensive description of the company, including its history, business model, and key products or services; discusses the company’s organizational structure, management team, and corporate governance; analyzes the company’s mission, vision, and strategic objectives.
Industry Analysis Provides an in-depth analysis of the industry in which the company operates; includes information on market size, growth trends, and key drivers; discusses the regulatory environment, technological advancements, and other external factors affecting the industry; analyzes the industry’s competitive dynamics and the company’s position within the industry.
Competitive Landscape Identifies the company’s main competitors and their market share; compares the company’s products, services, and pricing strategies with those of its competitors; analyzes the strengths and weaknesses of the company and its competitors; discusses potential new entrants and substitutes that could disrupt the competitive landscape.
Business Segments Provides a detailed analysis of the company’s various business segments or divisions; discusses the financial performance, growth prospects, and challenges of each segment; analyzes the contribution of each segment to the company’s overall revenue and profitability.
Growth Strategies Discusses the company’s strategies for driving future growth, such as organic growth initiatives, product innovations, and geographic expansions; analyzes the company’s mergers and acquisitions (M&A) strategy and potential targets; examines the company’s investments in research and development (R&D) and marketing.
Financial Analysis
Income Statement Analysis Analyzes the company’s revenue, expenses, and profitability.
Balance Sheet Analysis Examines the company’s assets, liabilities, and shareholders’ equity.
Cash Flow Analysis Analyzes the company’s cash inflows and outflows to evaluate liquidity.
Financial Ratios Discusses key ratios like profitability, liquidity, and solvency ratios.
Strategic Outlook
Investment Thesis Summarizes the key reasons for investing (or not investing) in the company’s shares; discusses the potential catalysts and risks that could impact the company’s valuation and stock price performance; provides a target price or price range for the company’s shares based on the valuation analyses and investment thesis.
Valuation Estimates the intrinsic value of the company’s shares using various valuation methodologies, such as discounted cash flow (DCF) analysis, relative valuation multiples, and sum-of-the-parts analysis; compares the company’s valuation with that of its peers and historical benchmarks; discusses the key assumptions and sensitivities underlying the valuation analyses.
Risk and Governance
Risk Factors Identifies and analyzes the key risks facing the company, such as market risks, operational risks, financial risks, and legal/regulatory risks; discusses the potential impact of these risks on the company’s financial performance and growth prospects; examines the company’s risk management strategies and mitigation measures.
Management and Governance Provides an overview of the company’s management team, including their experience, expertise, and track record; analyzes the company’s corporate governance practices, such as board composition, executive compensation, and shareholder rights; discusses the company’s succession planning and key person risks.
ESG Analyzes the company’s performance and initiatives related to environmental sustainability, social responsibility, and corporate governance; discusses the potential impact of ESG factors on the company’s reputation, risk profile, and financial performance; examines the company’s compliance with ESG regulations and industry standards.
Additional Content
Appendices and Disclosures Includes additional supporting materials, such as financial statements, ratio calculations, and detailed segment data; provides important disclosures, such as the analyst’s rating system, potential conflicts of interest, and disclaimers; discusses the limitations and uncertainties of the analysis and the need for further due diligence by investors.
None of the Above Covers any content that does not fall into the specified topics.
Table A4: Sentiment Portfolio Statistics
This table reports the value-weighted decile portfolio performances sorted by the average of four sentiment measures of the past 12 months’ analyst reports. Contemporaneous Return denotes two-day DGTW characteristic-adjusted buy-and-hold abnormal returns following report release dates. Recommendation Revision denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same stock. Headline Sentiment represents the sentiment score of report headlines measured using a fine-tuned BERT model. Body Sentiment represents the average sentiment score of report body content measured using a fine-tuned BERT model. For each portfolio, I report several performance measures: excess return mean and standard deviation, Sharpe ratio, $\alpha$ relative to Fama and French (2015) five factors with momentum (six factors in total), and the corresponding t-statistics for $\alpha$. Portfolios are rebalanced monthly. The ‘H-L’ row represents a long-short strategy that takes a long position in the highest decile (High) and a short position in the lowest decile (Low). The sample period extends from January 2005 to June 2024.
Contemporaneous Return Recommendation Revision
Mean SD SR $\alpha$ $t_{\alpha}$ Mean SD SR $\alpha$ $t_{\alpha}$
Low (L) 1.10 6.66 0.57 0.26 1.42 0.84 4.70 0.62 0.01 0.11
2 0.93 5.99 0.54 0.14 0.86 0.87 4.81 0.62 0.09 0.95
3 0.71 5.19 0.47 -0.01 -0.10 0.88 4.61 0.66 0.10 0.75
4 0.82 4.45 0.64 0.11 1.21 0.46 4.30 0.37 -0.33 -3.63
5 0.80 4.53 0.61 -0.08 -0.93 0.54 4.90 0.38 -0.20 -1.73
6 0.79 4.41 0.62 0.00 0.03 0.69 5.31 0.45 -0.12 -1.07
7 0.84 4.56 0.64 0.00 0.05 0.81 5.61 0.50 -0.04 -0.29
8 0.94 4.78 0.68 0.08 0.75 1.02 4.90 0.72 0.15 1.32
9 1.12 5.43 0.72 0.27 2.02 0.77 4.54 0.59 0.01 0.08
High (H) 0.78 6.06 0.44 -0.10 -0.72 1.11 4.73 0.81 0.25 1.94
H-L -0.33 4.20 -0.27 -0.36 -1.57 0.27 2.68 0.35 0.24 1.56
Headline Sentiment Body Sentiment
Mean SD SR $\alpha$ $t_{\alpha}$ Mean SD SR $\alpha$ $t_{\alpha}$
Low (L) 0.48 5.57 0.30 -0.23 -2.02 0.98 5.75 0.59 0.22 1.25
2 0.63 5.60 0.39 -0.09 -0.97 0.48 5.47 0.30 -0.17 -1.83
3 0.81 5.16 0.55 0.07 0.60 0.84 5.18 0.56 0.14 1.41
4 0.98 4.58 0.74 0.25 2.65 0.65 4.80 0.47 -0.09 -0.88
5 1.02 4.57 0.78 0.18 1.84 1.02 4.96 0.71 0.18 1.63
6 0.80 4.55 0.61 -0.02 -0.17 0.87 4.71 0.64 0.05 0.41
7 0.78 4.77 0.57 -0.08 -0.97 0.81 4.75 0.59 -0.10 -1.35
8 0.67 4.55 0.51 -0.19 -1.68 0.82 4.59 0.62 -0.03 -0.31
9 0.85 4.93 0.60 -0.03 -0.19 0.90 4.50 0.70 0.03 0.29
High (H) 0.74 5.11 0.50 -0.15 -1.50 0.82 4.56 0.63 -0.01 -0.16
H-L 0.27 3.31 0.28 0.08 0.58 -0.16 4.17 -0.13 -0.24 -0.97